OpenStudy (anonymous):

the perimeter of a square is 160 feet. what is the area of the square?

OpenStudy (calculusxy):

WHat is 160/4=

OpenStudy (anonymous):

If the perimetr, the sum of all 4 equal sides is 160, then each side is 40. That is 160/4 = 40. So the area is s^2 or 40 x 40 = 1,600 square feet.

OpenStudy (calculusxy):

We know that a sqare has four equal sides.